# MULAN-WC: Multi-Robot Localization Uncertainty-aware Active NeRF with Wireless Coordination

### v1

@@ Abstract
This paper presents MULAN-WC, a novel multi-robot 3D reconstruction framework that leverages wireless signal-based coordination between robots and Neural Radiance Fields (NeRF). Our approach addresses key challenges in multi-robot 3D reconstruction, including inter-robot pose estimation, localization uncertainty quantification, and active best-next-view selection. We introduce a method for using wireless Angle-of-Arrival (AoA) and ranging measurements to estimate relative poses between robots, as well as quantifying and incorporating the uncertainty embedded in the wireless localization of these pose estimates into the NeRF training loss to mitigate the impact of inaccurate camera poses. Furthermore, we propose an active view selection approach that accounts for robot pose uncertainty when determining the next-best views to improve the 3D reconstruction, enabling faster convergence through intelligent view selection. Extensive experiments on both synthetic and real-world datasets demonstrate the effectiveness of our framework in theory and in practice. Leveraging wireless coordination and localization uncertainty-aware training, MULAN-WC can achieve high-quality 3d reconstruction which is close to applying the ground truth camera poses. Furthermore, the quantification of the information gain from a novel view enables consistent rendering quality improvement with incrementally captured images by commending the robot the novel view position. Our hardware experiments showcase the practicality of deploying MULAN-WC to real robotic systems.
@@ I Introduction
Vision-based 3D reconstruction in previously unseen environments is pivotal in a broad spectrum of robotics applications, ranging from autonomous navigation[1], mapping and localization[2] to scene understanding[3]. The conventional process typically involves: 1) collecting multi-modal sensory information from onboard sensors such as RGB-D cameras and inertial measurement units, 2) extracting geometric features to compute relative pose information, and 3) applying pose graph optimization to produce a 3D environment representation using geometrically constrained spatial feature information[4, 5, 6]. Scaling up this capability to a fleet of robots could enable better coverage and faster exploration in large-scale environments. Nevertheless, it is nontrivial to scale up conventional methods to a fleet of robots. This is due to challenges in effectively obtaining relative poses between robots that are needed to align inter-robot frames and form a global understanding of the scene. Furthermore, another problem in deploying a multi-robot system is the question of how to actively command the robot to acquire visual information so as to maximize the information gain in the 3D reconstruction [7, 8]. Active image acquisition is even more critical in a multi-robot setting to fully leverage the advantages of the fleet over a single robot. To address these challenges, we introduce a multi-robot collaborative framework utilizing Neural Radiance Fields (NeRF) for reconstruction, and using on-board wireless signal-based coordination to provide relative positional information between robots.
@@
Firstly, to address the need for photometrically and geometrically accurate 3d reconstruction, a large number of works in NeRF[9, 10, 11] offer a revolutionary technique in synthesizing photorealistic 3D implicit representations from sparse 2D images. This is attributed to NeRF’s unique capability to model the volumetric density and color of light in a scene, enabling highly detailed and accurate reconstructions from diverse viewpoints. A crucial input that enables real-time NeRF in [11] is the camera pose corresponding to each image in the same frame. Using optical flow based feature tracking described in [12], the relative camera poses can be computed in real-time and fed to the NeRF training. However, acquiring accurate relative positional information in a multi-robot system is nontrivial, especially in the absence of global localization systems like GNSS or motion capture systems. In the traditional multi-robot coordination or localization approach, multi-robot SLAM is often dependent on the alignment of individual maps and subsequent pose estimation from overlapped appearance-based feature observations [5, 13], namely loop closure. However, inter-robot loop closure brings significant complications and computational overhead [14, 5]. To satisfy the need for relative positional information, we instead use phased array-based wireless sensing between robots, building upon our previous work in[15]. Here, the off-the-shelf WiFi chip, which is native for communication on most robotics platforms, can be used to obtain inter-robot positional information independent of appearance-based environmental features. This positional information thus can be utilized to compute the relative translation between any pair of robots and thus their image frames. Inspired by other works in incorporating depth information from the SLAM pipeline[12] where depth information is used as supervision, we also design our training loss to be aware of which regions of data are more certain than others based on the uncertainty in wireless sensing. Like other perception modalities, wireless sensing also encodes probabilistic perception due to environmental and hardware noise. This work develops a method to quantify the wireless localization uncertainty based on Angle of Arrival (AoA) profile reconstruction and correlation with the received AoA profile. This ensures that the training loss is informed about data regions with higher localization uncertainty, which brings in more accurate visual information. Integrating this quantified uncertainty information into the NeRF training process allows us to bias the training loss, enhancing the accuracy and reliability of the 3D NeRF-generated reconstructions.
@@
Moreover, a multi-robot system offers advantages beyond merely improving the efficiency of scene coverage from different viewpoints[16]; we can also naturally enable active image acquisition by determining the most informative next view for the NeRF model and controlling robots to acquire these additional images. Most works in applying NeRF to 3D reconstruction only passively process the images given by the perception pipeline. In resource-constrained large-scale deployment, it is beneficial to actively plan robots to acquire the most informative next image. There are some works [17, 18] that acquire images that can maximally cover the scene of interest, leading to higher information gain or better quality of reconstruction. However, to the best of our knowledge, this is the first time that active information acquisition has been applied to a coordinated multi-robot system for NeRF-based 3D reconstruction. The work in [17] proposes a promising approach to evaluate the potential information gained from a novel view by quantifying the reduction of the variance for the rendering. However, this quantification doesn’t consider the localization uncertainty of the camera, which is particularly common in multi-robot or multiple-camera setups, making it inefficient while dealing with localization uncertainty across robots that use wireless coordination. We address this specifically, by considering the inter-robot camera pose uncertainty in the characterization of the color posterior in 3D space. Our work integrates localization uncertainty quantification into the evaluation of novel-view information gain by deriving the reduction of the variance. Subsequently, we can direct the robots to actively capture images from a set of feasible next positions, for the team of robots to achieve the highest information gain in the NeRF model.
@@
  • Extensive Experiments on hardware robots: We conducted experiments on our customized hardware robot demonstrating that our framework does not only effectively achieve the same quality of rendering faster and higher quality, but also can actively command the robot to unvisited place by reducing the variance of rendering.
@@ II Problem formulation
In this section, we briefly review some background knowledge of the NeRF and introduce the wireless coordination from our previous work [15]as a basis for our approach.
@@ II-A NeRF formulation
NeRF implicitly represents a scene using a fully connected neural network. In the ideal propagation ray-tracing model, the scene is modeled as a continuous function that maps any viewing angle in 5D input coordinates to a color $c(r,g,b)$ and a volume density $\sigma$. The 5D input coordinate $D$ consists of the position in Cartesian coordinates $(x,y,z)$ and the viewing angle $(\theta,\phi)$. NeRF renders the color of the sampling ray passing through the environment with classical volume rendering. Suppose we sample a ray from a position $\boldsymbol{o}$. The points along the ray can be parameterized by
@@ II-B Collaborative NeRF
To achieve 3D reconstruction with more than one robot, one of the fundamental requirements is having a common frame of reference or known camera extrinsic or relative transformation between cameras even if the data is collected from different robots from different views. Instead of being given a set of poses $\mathcal{T}$ in the same frame of reference, we instead focus on the problem of having the sets of poses from all robots in the team in $\mathcal{T}_{\alpha}$, $\mathcal{T}_{\beta}$ and so on. Without loss of generality, we only focus on the observation from two robots $\alpha$ and $\beta$. In order to align a pose $T_{k}^{\alpha}$ of robot $\alpha$ at local time k and another pose $T_{p}^{\beta}$ of robot $\beta$ at local time p, we need to obtain the inter-robot camera extrinsic $T_{k_{\alpha}}^{p\beta}=(t_{k_{\alpha}}^{p\beta},\theta_{k_{\alpha}}^{p\beta})$.
@@ II-C Wireless Coordination
In our previous work [14], we extract the Angle of Arrival(AoA) information between any two robots by measuring the phase difference in the Wi-Fi channel. Suppose we have two robots $\alpha$ and $\beta$ in communicating range at time $t$ and their poses in local frame $T_{\alpha}$ and $T_{\beta}$. We can measure relative position between two robots using ranging from the ultrawideband (UWB) as well as Angle-of-Arrival (AoA) from our SAR-based framework output [15] with a probability density function of ranging and AoA annotated by $f_{uwb}(d|T^{\alpha},T^{\beta})$ and $f_{aoa}(\phi|T^{\alpha},T^{\beta})$ respectively, are defined as:
@@
where $c_{1}=\frac{1}{\sqrt{2\pi\sigma_{k,p}^{2}}}$, $c_{2}=\frac{1}{2\pi I_{0}(\kappa_{k,p})}$. Here, $\sigma_{k,p}^{2}$ and $t_{k\alpha}^{p\beta}$ the variance and mean of the distance measurement; and $\kappa^{2}_{k,p}$ and $\theta_{k\alpha}^{p\beta}$ are the concentration parameter computed as the inverse of the AoA variance and the mean of the AoA distribution.
@@ III Approach
In this section, we present our multi-robot NeRF framework that addresses the challenges of inter-robot pose localization, uncertainty quantification, and active best-next-view finding. Our approach leverages wireless signals, specifically Angle of Arrival (AoA) and ranging measurements, to estimate the relative poses between robots. We develop a novel method to quantify the uncertainty of AoA estimates by reconstructing the AoA profile and correlating it with the received AoA profile. This uncertainty quantification is then integrated into the NeRF training process to mitigate the impact of inaccurate poses on the reconstruction quality. Furthermore, we propose an active view-finding approach that accounts for the position uncertainty of the robots when selecting the most informative views for NeRF training. By incorporating localization uncertainty into the novel view selection process, our framework can more accurately determine the best next views for each robot, even in the presence of pose uncertainty arising from wireless coordination.
@@ III-A Inter-robot Pose Localization and Uncertainty Quantification
As described in Section II-B, accurate and fast NeRF reconstruction relies heavily on the known transformation between cameras on different robots. In a multi-robot setup, we propose using wireless signals to obtain accurate inter-robot poses by leveraging Angle of Arrival (AoA) and ranging measurements. Suppose we have a pose $T_{k}^{\alpha}$ in SE(3) of robot $\alpha$ at local time k and another pose $T_{p}^{\beta}$ of robot $\beta$ at local time p. We then can obtain a wireless measurement between the two robots using onboard WiFi and UWB which can be annotated by a tuple $(t_{k_{\alpha}}^{p\beta},\theta_{k_{\alpha}}^{p\beta})$. If we aim to use robot $\alpha$’s frame as the global frame, then the extrinsic or the rigid transformation of robot $\beta$’s camera pose can be represented by $T_{kp}^{\alpha\beta}\oplus T_{p}^{\beta}$, where the annotation $\oplus$ denotes rigid transformation. However, the accuracy of the resulting pose estimates is subject to the uncertainties in the AoA and ranging measurements as described in Eq 2 and Eq 3. To mitigate the impact of inaccurate poses on the NeRF training process, we propose applying a weight to each training example based on the uncertainty of the associated robot pose’s AoA and ranging measurements from the other robots.
Quantifying the uncertainty of AoA estimates is particularly challenging, since there is a lack of standard error quantification methods applicable from previous works. To address this, we develop a novel approach to quantify the uncertainty of AoA estimates by reconstructing the AoA profile and correlating it with the received AoA profile. We propose the AoA uncertainty quantification methodology with the following outline. The receiving robot $\alpha$’s wireless channel and pose $<\overline{h}_{\alpha\beta}(t),\overline{T}_{\alpha}(t)>$ are collected over some set of signal packet timestamps $t=t_{k},\dots,t_{l}$ in robot $\alpha^{\prime}s$ timeframe from a transmitting robot $\beta$. We want to measure AoA of robot $\beta$ at robot $\alpha$’s local time $t_{k}$. Following [20], the poses of robot $\alpha$ trace out a virtual multi-antenna array that enables AoA measurement. First, we apply multiple signal classification (MUSIC) to the measured channels over the timestamps to obtain a measured AoA profile $\overline{F}(\phi,\theta)$ and AoA measurement $(\overline{\phi},\overline{\theta})$. Then, we reconstruct the wireless channel phases in $h_{\alpha\beta}^{\prime}(t)$ at each timestamp taking the transmitter to be at the measured AoA, apply MUSIC to obtain the reconstructed AoA profile $F^{\prime}(\phi,\theta)$, and calculate the AoA uncertainty by comparing two profiles to measure the similarity of the profiles near $(\overline{\phi},\overline{\theta})$.
Figure 1: Simulation of the AoA variance methodology. (a) The Measured AoA profile $\overline{F}_{\alpha\beta}(\phi,\theta)$ from a simulated measured wireless channel $\overline{h}_{\alpha\beta}$ with $0.7$ radians standard deviation injected phase noise. (b) The reconstructed AoA profile $F_{\alpha\beta}^{\prime}(\phi,\theta)$ from the reconstructed channel $h_{\alpha\beta}^{\prime}$, with $0.5$ radians standard deviation phase noise for tolerance. Note both tallest peaks align at $(\overline{\phi}=45.6^{\circ},\overline{\theta}=90^{\circ})$. (c) and (d) are top views of (a) and (b), respectively.
Specifically, the ideal channel of robot $\alpha$ from robot $\beta$ is
$h_{\alpha\beta}(t)=\frac{1}{d_{\alpha\beta}}exp\left(\frac{-2\pi\sqrt{-1}}{\lambda}d_{\alpha\beta}(t)\right)$ (4)
@@
and robot $\alpha$ collects its local pose information, particularly the displacement distance, azimuth, and zenith of robot $i$ from the center of its frame, over $t=t_{k},\dots,t_{l}$. Suppose robot $i$ receives a measured channel $\overline{h}_{\alpha\beta}(t)$. Then the MUSIC AoA profile is constructed as in [20] to be $\overline{F}_{\alpha\beta}(\phi,\theta)$, and the measured AoA is taken to be $(\overline{\phi},\overline{\theta})=\arg\max_{(\phi,\theta)}\left\{\overline{F}_{\alpha\beta}(\phi,\theta)\right\}$ at the tallest peak in the profile.
@@
From here, we reconstruct the channel over $t=t_{k},\dots,t_{l}$ using $(\overline{\phi},\overline{\theta})$ and the pose information $\overline{T}_{\alpha}(t_{k}),\dots,\overline{T}_{\alpha}(t_{l})$:
@@
$h_{\alpha\beta}^{\prime}(t)=exp\left(\frac{-2\pi\sqrt{-1}}{\lambda}f_{\alpha}(\overline{T}_{\alpha}(t),\overline{\phi},\overline{\theta})+\sqrt{-1}\nu(t)\right)$ (5)
@@
where $f_{\alpha}$ is the displacement of robot $i$ projected along the measured AoA direction all in the local frame, relative to the first observation at $t_{k}$, and $\nu(t)$ is a zero-mean real random variable that injects Gaussian phase noise into each element of the channel to add a small realistic amount of error tolerance in the measured profile. The same MUSIC algorithm is then run to obtain $F_{\alpha\beta}^{\prime}(\phi,\theta)$ and its tallest peak at $(\phi^{\prime},\theta^{\prime})$. Since the reconstructed channel is based on an AoA from $(\overline{\phi},\overline{\theta})$, making the noise $\nu(t)$ small ensures that the reconstruction has $(\phi^{\prime},\theta^{\prime})=(\overline{\phi},\overline{\theta})$, where the $\phi$ and $\theta$ sample spaces are discretized during MUSIC computation.
@@
The AoA uncertainty is then calculated as follows. Both profiles are cropped to a rectangle $R$ with a small radius $\Delta\phi=5^{\circ}$ and $\Delta\theta=5^{\circ}$ around $(\overline{\phi},\overline{\theta})$ for comparison near the measured peak, to ignore distant multipath components. Then we define the AoA uncertainty $\kappa_{k,p}$ via
@@
This measures how concentrated the received profile is around the $(\overline{\phi},\overline{\theta})$, such that a lower AoA uncertainty $\kappa_{k,p}$ corresponds to a more reliable AoA measurement. This AoA variance can be used in Equation 3 to understand the variability in the pose estimates, as described in the next section.

### v2

@@ Abstract
This paper presents MULAN-WC, a novel multi-robot 3D reconstruction framework that leverages wireless signal-based coordination between robots and Neural Radiance Fields (NeRF). Our approach addresses key challenges in multi-robot 3D reconstruction, including inter-robot pose estimation, localization uncertainty quantification, and active best-next-view selection. We introduce a method for using wireless Angle-of-Arrival (AoA) and ranging measurements to estimate relative poses between robots, as well as quantifying and incorporating the uncertainty embedded in the wireless localization of these pose estimates into the NeRF training loss to mitigate the impact of inaccurate camera poses. Furthermore, we propose an active view selection approach that accounts for robot pose uncertainty when determining the best-next-views to improve the 3D reconstruction, enabling faster convergence through intelligent view selection. Extensive experiments on both synthetic and real-world datasets demonstrate the effectiveness of our framework in theory and in practice. Leveraging wireless coordination and localization uncertainty-aware training, MULAN-WC can achieve high-quality 3D reconstruction that is close to applying the ground truth camera poses. Furthermore, the quantification of the information gain from a novel view enables consistent rendering quality improvement with incrementally captured images by commanding the robot to the novel view position. Our hardware experiments showcase the practicality of deploying MULAN-WC to real robotic systems.
@@ I Introduction
Vision-based 3D reconstruction in previously unseen environments is pivotal in a broad spectrum of robotics applications, ranging from autonomous navigation [1] to mapping and localization [2] to scene understanding [3]. The conventional process typically involves: 1) collecting multi-modal sensory information from onboard sensors such as RGB-D cameras and inertial measurement units, 2) extracting geometric features to compute relative pose information, and 3) applying pose graph optimization to produce a 3D environment representation using geometrically constrained spatial feature information [4, 5, 6]. Scaling up this capability to a fleet of robots could enable better coverage and faster exploration in large-scale environments. Nevertheless, it is nontrivial to scale up conventional methods to a fleet of robots. This is due to challenges in effectively obtaining relative poses between robots, which are needed to align inter-robot frames and form a global understanding of the scene. Another problem lies in how to actively command the robot to acquire visual information so as to maximize the information gain in the 3D reconstruction [7, 8]. Active image acquisition is even more critical in a multi-robot setting to fully leverage the advantages of the fleet over a single robot. To address these challenges, we introduce a multi-robot collaborative framework utilizing Neural Radiance Fields (NeRF) for reconstruction, and using on-board wireless signal-based coordination to provide relative positional information between robots.
@@
Firstly, to address the need for photometrically and geometrically accurate 3D reconstruction, a large number of works in NeRF [9, 10, 11] offer a revolutionary technique in synthesizing photorealistic 3D implicit representations from sparse 2D images. This is attributed to NeRF’s unique capability to model the volumetric density and color of light in a scene, enabling highly detailed and accurate reconstructions from diverse viewpoints. A crucial input that enables real-time NeRF in [11] is the camera pose corresponding to each image in the same frame. Using optical flow-based feature tracking described in [12], the relative camera poses can be computed in real-time and fed into the NeRF training. However, acquiring accurate relative positional information in a multi-robot system is nontrivial, especially in the absence of global localization systems like GNSS or motion capture systems. In the traditional multi-robot coordination or localization approach, multi-robot SLAM is often dependent on the alignment of individual maps and subsequent pose estimation from overlapped appearance-based feature observations [5, 13], namely loop closure. However, inter-robot loop closure brings significant complications and computational overhead [14, 5, 15]. To satisfy the need for relative positional information, we instead use phased array-based wireless sensing between robots, building upon our previous work in [16, 17]. Here, the off-the-shelf WiFi chip, which is native for communication on most robotics platforms, can be used to obtain inter-robot positional information independent of appearance-based environmental features. This positional information thus can be utilized to compute the relative translation between any pair of robots and thus their image frames. Inspired by other works in incorporating depth information from the SLAM pipeline [12] where depth information is used as supervision, we also design our training loss to be aware of which regions of data are more certain than others based on the uncertainty in wireless sensing. Like other perception modalities, wireless sensing also encodes probabilistic perception due to environmental and hardware noise. This work develops a method to quantify the wireless localization uncertainty based on Angle of Arrival (AoA) profile reconstruction and correlation with the received AoA profile. Integrating this quantified uncertainty information into the NeRF training process allows us to bias the training loss and ensure that the training loss is informed about data regions with higher localization uncertainty, enhancing the accuracy and reliability of the 3D NeRF-generated reconstructions.
@@
A multi-robot system offers advantages beyond merely improving the efficiency of scene coverage from different viewpoints [18]; we can also enable active image acquisition by determining the most informative next view for the NeRF model and controlling robots to acquire these additional images. Most works in applying NeRF to 3D reconstruction only passively process the images given by the perception pipeline. In resource-constrained large-scale deployment, it is beneficial to actively plan robots to acquire the most informative next image. Some works [19, 20] acquire images that can maximally cover the scene of interest, leading to higher information gain or better quality of reconstruction. However, to the best of our knowledge, this is the first time that active information acquisition has been applied to a coordinated multi-robot system for NeRF-based 3D reconstruction. The work in [19] proposes a promising approach to evaluate the potential information gained from a novel view by quantifying the reduction of the variance for the rendering. However, this quantification does not consider the localization uncertainty of the camera, which is particularly necessary in multi-robot or multiple-camera setups, making it inefficient when dealing with localization uncertainty across robots that use wireless coordination. We address this specifically by considering the inter-robot camera pose uncertainty in the characterization of the color posterior in 3D space. Our work integrates localization uncertainty quantification into the evaluation of novel-view information gain by deriving the reduction of the variance. Subsequently, we can direct the robots to actively capture the best images from a set of feasible next positions, for the team of robots to achieve the highest information gain in the NeRF model.
@@
  • Extensive hardware experiments: We conducted experiments on our customized hardware robot demonstrating that our framework does not only effectively achieve the same quality of rendering faster and higher converged quality, but also can actively command the robot to an unvisited place that reduces the variance of rendering.
@@ II Problem formulation
In this section, we briefly review some background knowledge of NeRF and introduce the wireless coordination from our previous work [17] as a basis for our approach.
@@ II-A NeRF Formulation
NeRF implicitly represents a scene using a fully connected neural network. In the ideal propagation ray-tracing model, the scene is modeled as a continuous function that maps any viewing angle $D$ of 5D input coordinates, consisting of the position in Cartesian coordinates $(x,y,z)$ and the viewing angle $(\theta,\phi)$, to a color $c(r,g,b)$ and a volume density $\sigma$. NeRF renders the color of the sampling ray passing through the environment with classical volume rendering. Suppose we sample a ray from a position $\boldsymbol{o}$ in direction $\boldsymbol{d}$. The points along the ray can be parameterized by
@@
where $\bar{\mathcal{C}}(r_{i})$ is the captured color from images.
@@ II-B Collaborative NeRF
To achieve 3D reconstruction with more than one robot, one of the fundamental requirements is having a common frame of reference from a known camera extrinsic or relative transformation between cameras even if the data is collected from different robots from different views. Instead of being given a set of poses $\mathcal{T}$ in the same frame of reference, we instead focus on the problem of having the sets of poses from all the robots $\alpha,\beta,\dots$ in the team in $\mathcal{T}_{\alpha}$, $\mathcal{T}_{\beta},\dots$. Without loss of generality, we only focus on the observation from two robots $\alpha$ and $\beta$. In order to align a pose $T_{k}^{\alpha}$ of robot $\alpha$ at local time k and another pose $T_{p}^{\beta}$ of robot $\beta$ at local time p, we need to obtain the inter-robot camera extrinsic $T_{k_{\alpha}}^{p\beta}=(t_{k_{\alpha}}^{p\beta},\theta_{k_{\alpha}}^{p\beta})$, which is the distance and Angle of Arrival (AoA) between two cameras on different robots.
@@ II-C Wireless Coordination
In our previous work [14], we extract the AoA information between any two robots by measuring the phase difference in the Wi-Fi channel. Suppose we have two robots $\alpha$ and $\beta$ in communicating range at time $t$ and their poses $T_{\alpha}$ and $T_{\beta}$ in local frames. We can measure relative position between two robots using ranging from the ultra-wideband (UWB) as well as AoA from our SAR-based framework output [17] with a probability density function of ranging and AoA annotated by $f_{uwb}(d|T^{\alpha},T^{\beta})$ and $f_{aoa}(\phi|T^{\alpha},T^{\beta})$ respectively, defined as:
@@
where $c_{1}=\frac{1}{\sqrt{2\pi\sigma_{k,p}^{2}}}$ and $c_{2}=\frac{1}{2\pi I_{0}(\kappa_{k,p})}$. Here, $\sigma_{k,p}^{2}$ and $t_{k\alpha}^{p\beta}$ are the variance and mean of the distance measurement; and $\kappa^{2}_{k,p}$ and $\theta_{k\alpha}^{p\beta}$ are the concentration parameters computed as the inverse of the AoA variance and the mean of the AoA distribution.
@@ III Approach
In this section, we present our multi-robot NeRF framework that addresses the challenges of inter-robot pose localization, uncertainty quantification, and active best-next-view finding. Our approach leverages wireless signals, specifically ranging and Angle of Arrival (AoA) measurements, to estimate the relative poses between robots. We develop a novel method to quantify the uncertainty of AoA estimates by reconstructing the AoA profile and correlating it with the received AoA profile. This uncertainty quantification is then integrated into the NeRF training process to mitigate the impact of inaccurate poses on the reconstruction quality. Furthermore, we propose an active view-finding approach that accounts for the position uncertainty of the robots when selecting the most informative views for NeRF training. By incorporating localization uncertainty into the novel view selection process, our framework can more accurately determine the best next views for each robot, even in the presence of pose uncertainty arising from wireless coordination.
@@ III-A Inter-robot Pose Localization and Uncertainty Quantification
As described in Section II-B, accurate multi-robot NeRF reconstruction relies on obtaining the transformation between cameras on different robots. In a multi-robot setup, we propose using wireless signals leveraging UWB ranging and WiFi AoA measurements to obtain accurate inter-robot poses. Suppose we have a pose $T_{k}^{\alpha}$ in SE(3) of robot $\alpha$ at local time k and another pose $T_{p}^{\beta}$ of robot $\beta$ at local time p. We then can obtain a wireless measurement of range and AoA between the two robots using onboard UWB and WiFi annotated by a tuple $(t_{k_{\alpha}}^{p\beta},\theta_{k_{\alpha}}^{p\beta})$. If we aim to use robot $\alpha$’s frame as the global frame, then the extrinsic or the rigid transformation of robot $\beta$’s camera pose can be represented by $T_{kp}^{\alpha\beta}\oplus T_{p}^{\beta}$, where the annotation $\oplus$ denotes rigid transformation. However, the accuracy of the resulting pose estimate is subject to the uncertainties in the ranging and AoA measurements as described in Eq 2 and Eq 3. To mitigate the impact of inaccurate poses on the NeRF training process, we propose applying a weight to each training example based on the uncertainty of the associated robot pose’s ranging and AoA measurements from the other robots.
Quantifying the uncertainty of AoA estimates is particularly challenging, since there is a lack of standard error quantification methods applicable from previous works. To address this, we propose a novel approach as follows. The ideal channel on wavelength $\lambda$ at robot $\alpha$ from robot $\beta$ over distance $d_{\alpha\beta}(t)$ is
$h_{\alpha\beta}(t)=\frac{1}{d_{\alpha\beta}(t)}exp\left(\frac{-2\pi\sqrt{-1}}{\lambda}d_{\alpha\beta}(t)\right)$ (4)
@@
Suppose over $t=t_{k},\dots,t_{l}$, robot $\alpha$ receives the measured channel $\overline{h}_{\alpha\beta}(t)$ and also collects its local pose information $\overline{T}_{\alpha}(t)$ containing the displacement distance, azimuth, and zenith of robot $\alpha$ from the center of its frame. Then the measured AoA profile is constructed as in [16] to be $\overline{F}_{\alpha\beta}(\phi,\theta)$ over a sample space in $(\phi,\theta)$ and the measured AoA is chosen as $(\overline{\phi},\overline{\theta})=\arg\max_{(\phi,\theta)}\left\{\overline{F}_{\alpha\beta}(\phi,\theta)\right\}$ at the tallest peak.
@@
Now, given $(\overline{\phi},\overline{\theta})$ and pose information $\overline{T}_{\alpha}(t_{k}),\dots,\overline{T}_{\alpha}(t_{l})$, we reconstruct the channel over $t=t_{k},\dots,t_{l}$ as
@@
$h_{\alpha\beta}^{\prime}(t)=exp\left(\frac{-2\pi\sqrt{-1}}{\lambda}f_{\alpha}(\overline{T}_{\alpha}(t),\overline{\phi},\overline{\theta})+\nu(t)\sqrt{-1}\right)$ (5)
@@
where $f_{\alpha}$ is the displacement of robot $\alpha$ projected along the measured AoA direction in the local frame, relative to the first observation at $t_{k}$, and $\nu(t)$ is a zero-mean real random variable that injects Gaussian phase noise into each element of the channel to add a small realistic amount of error tolerance in the measured profile. The same AoA algorithm is run to obtain the reconstructed profile $F_{\alpha\beta}^{\prime}(\phi,\theta)$ and its tallest peak $(\phi^{\prime},\theta^{\prime})$. Since $F_{\alpha\beta}^{\prime}(\phi,\theta)$ is constructed from $(\overline{\phi},\overline{\theta})$, making the noise $\nu(t)$ small ensures that the reconstruction has $(\phi^{\prime},\theta^{\prime})=(\overline{\phi},\overline{\theta})$ when the $\phi$ and $\theta$ sample spaces are discretized during profile computation. Figure 1 simulates the profiles for illustration.
Figure 1: Simulation of the AoA variance methodology. (a) Measured AoA profile $\overline{F}_{\alpha\beta}(\phi,\theta)$ from a simulated measured wireless channel $\overline{h}_{\alpha\beta}$ with $0.7$ radians standard deviation injected phase noise. (b) Reconstructed profile $F_{\alpha\beta}^{\prime}(\phi,\theta)$ from reconstructed channel $h_{\alpha\beta}^{\prime}$, with $0.5$ radians standard deviation phase noise for tolerance. Both tallest peaks align at $(\overline{\phi}=45.6^{\circ},\overline{\theta}=90^{\circ})$. (c) and (d) are respective top views.
@@
The AoA uncertainty is then calculated as follows. Both profiles are cropped to a small rectangle $R$ with $\Delta\phi=\Delta\theta=10^{\circ}$ around the measured AoA $(\overline{\phi},\overline{\theta})$ to ignore distant multipath components. Then define AoA uncertainty $\kappa_{k,p}$ via
@@
This measures how concentrated the received profile is around its peak $(\overline{\phi},\overline{\theta})$, such that a lower AoA uncertainty $\kappa_{k,p}$ corresponds to a more reliable AoA measurement. This AoA variance can be used in Equation 3 to understand the variability in the pose estimates, as described in the next section.